3x2 + 2x = 7


x2 + 8x+9=-7

Answer by stanbon(75887)

3x2 + 2x = 7
3x^2+2x-7 = 0
x = [-2 +- sqrt(4-4*3*-7)]/6
x = [-2 +- sqrt(89)]/6
x = [-2 + sqrt(89)]/6
x = [-2 - sqrt(89)]/6
-------------------------
x^2 + 8x + 9 = -7
x^2+8x+16=0
(x+4)^2 = 0
x = -4 with multiplicity 2
========================
Cheers,
Stan H.